OpenStudy (anonymous):

2/3x+y=7 1/2x+3/4y =21/4 I know it has more then one solution i just need to know the formula to make it right it is written as t being x and y being some number times x divided by another number

OpenStudy (babyx3boo):

Would the formula be the y-int formular y=mx+b?

OpenStudy (babyx3boo):

formula*

OpenStudy (anonymous):

its set up as two systems you have to really find y as x can be any number as its infinite solutions

OpenStudy (mathmale):

JC: Your 2/3x+y=7 1/2x+3/4y =21/4 is, unfortunately, hard to read and thus is subject to misinterpretation. For example, do you mean (2/3)x + y = 7, or is that (2/3)x+y= (7 1/2)x+3, or what? It'd be so much clearer if you'd please draw the system of equations straight from the original; use the Draw utility, below. Or, use parentheses liberally and leave more room to separate one equation from the next. 

OpenStudy (mathmale):

Also, please be clearer about what the object of this problem is. Are you trying to solve a system of equations for x and y?